# Scaling Laws Revisited: Modeling the Role of Data Quality in Language Model Pretraining

### v1

@@ A.3 Proof of Proposition 1
By assumption (as in the Information Bottleneck principle [37]), the data-dependent loss scales as
@@ B.3.1 Paracrawl v8
In preparing our dataset for machine translation from English to German, we used a pretrained language detection model from FastText [24] lid.176.bin to filter the data as described in 5.1.
@@
We use a cleaned version of C4 by AllenAI [13] available on HuggingFace, they provide 5 splits, en, en.noclean, en.noblocklist, realnewslike, and multilingual (mC4). We use the en variant in our experiments, it contains 364,868,892 samples in the train split and 364,608 samples in the validation split.
@@ C.1 Additional Experimental Details
Table 5 shows other hyperparameters and choices made during our training experiments. We also notably use greedy dataset packing in NMT and make use of DataCollatorWithFlattening() from HuggingFace in CLM to train to easily use Flash Attention 2 [11]. Additionally we use weight decay of 0.01 in both cases.
@@ C.3 Parametric Loss Fitting with Huber Loss
Similar to [20] we minimize the following objective following a regularization introduced by [21]
@@
where $B$ = $\exp{(b)}$, $E^{+}$ = $\exp{(e)}$ and $\delta=10^{-3}$ using the L-BFGS-B solver available through SciPy minimize. We replicate the grid initialization of parameters as follows, $b\in[0,25,5]$, $e\in[0,2,0.5]$, $\beta\in[0.0,0.4,0.1]$ and $\gamma\in[0.0,0.4,0.1]$. We also add bounds $0\leq\beta\leq 1$ and $0\leq\gamma\leq 1$. We also report a secondary fit using SciPy inbuilt function curve fitting functions using Least Squares.

### v2

@@ A.3 Proof of Proposition 1
By assumption (as in the Information Bottleneck principle (Tishby et al., 1999)), the data-dependent loss scales as
@@ B.3.1 Paracrawl v8
In preparing our dataset for machine translation from English to German, we used a pretrained language detection model from FastText (Joulin et al., 2016) lid.176.bin to filter the data as described in 5.1.
@@
We use a cleaned version of C4 by AllenAI Dodge et al. (2021) available on HuggingFace, they provide 5 splits, en, en.noclean, en.noblocklist, realnewslike, and multilingual (mC4). We use the en variant in our experiments, it contains 364,868,892 samples in the train split and 364,608 samples in the validation split.
@@ C.1 Additional Experimental Details
Table 6 shows other hyperparameters and choices made during our training experiments. We also notably use greedy dataset packing in NMT and make use of DataCollatorWithFlattening() from HuggingFace in CLM to train to easily use Flash Attention 2 (Dao, 2023). Additionally we use weight decay of 0.01 in both cases.
@@ C.3 Parametric Loss Fitting with Huber Loss
Similar to Hoffmann et al. (2022) we minimize the following objective following a regularization introduced by Huber (1992)
@@
where $B$ = $\exp{(b)}$, $E$ = $\exp{(e)}$ and $\delta=10^{-3}$ using the L-BFGS-B solver available through SciPy minimize. We replicate the grid initialization of parameters as follows, $b\in[0,25,5]$, $e\in[0,2,0.5]$, $\beta\in[0.0,0.4,0.1]$ and $\gamma\in[0.0,0.4,0.1]$. We also add bounds $0\leq\beta\leq 1$ and $0\leq\gamma\leq 1$. We also report a secondary fit using SciPy inbuilt function curve fitting functions using Least Squares.
Appendix D Relationship to Other Definitions of Data Quality
In this section, we show that our quality aware scaling law recovers the scaling law in Chen et al. (2025) by assuming that the data deficiency $\Delta(\omega)$ for a dataset $\omega$ has the following form:
$\Delta(\omega)=\mu_{1}E+\mu_{2}\frac{1}{F}+\mu_{3}G$ (5)
where $E$ is the total noise, $F$ is the coverage scaled between 0 (minimum coverage) and $1$ (maximum coverage), and $G$ quantifies the amount of repeated data.
We assume that density as defined in Chen et al. (2025), which incorporates both coverage and repeated data, is quantified such that low density is associated with high quality data and high density is associated with lower quality data.
Let $\rho$ be the density of a dataset as defined in Chen et al. (2025), Equation 3, with the density assumed to satisfy $0\leq\rho\leq 1$. Here low $\rho$ close to 0 corresponds to higher quality data with fewer repetitions, and high $\rho$ close to 1 corresponds to low quality data with more repetitions, leading to sub-scaling behavior Chen et al. (2025).
We define the data dispersion to be
$\mbox{Dis}(\rho)=-\ln(1-\rho),$
so as $\rho$ approaches $1$ (low quality data in the sense of Chen et al. (2025)), $\mathrm{Dis}$ approaches $\infty$, and as $\rho$ approaches 0 (high quality data in the sense of Chen et al. (2025)), Dis approaches $1$.
In the equation above, we assume that the data dispersion $\mathrm{Dis}$ captures the coverage $1/F$ of the data and density $\rho$. We also assume that the $\mathrm{Dis}$ captures the amount of repeated data, although as we will see below, this can also be modeled in more precise ways. If we fix the noise $E$, then the data deficiency is
$\Delta(\omega)=\mathrm{constant}+\mu_{2}\mathrm{Dis}(\rho)$ (6)
and we recover the data quality scaling law in Chen et al. (2025). In particular, as the density $\rho$ approaches 0, the data quality factor $Q$ approaches 1, and as the density $\rho$ approaches $1$, the data quality degrades.
More generally, we can define a modified data dispersion
$\mathrm{Dis}^{\kappa}(\rho)=-\ln(1-\rho^{\kappa}).$
Turning towards the data quality scaling law in Chang et al. (2024), if instead of using the cluster based definition of density, we use the compression based definition of density, we get
$\mathrm{CR}(\omega)$ $=\mathrm{size}(\omega)/\mathrm{compressed\ size}(\omega)$
$\mathrm{DR}(\omega)$ $=\mathrm{CR}^{-1}(\omega)$
We define dispersion as:
$\mbox{Dis}(\rho)=-\ln(1-\mathrm{DR}(\omega)).$
With this definition of $\mathrm{Dis}(\omega)$, the rest of the argument is the same.
In a similar fashion, we can add a term $H=S(\omega)$ to the data deficiency $\Delta(\omega)$ to capture the increase in data deficiency due to the negative impact of synthetic data, which we can quantify with a term as follows
$\Delta(\omega)=\mu_{1}E+\mu_{2}\frac{1}{F}+\mu_{3}G+\mu_{4}H$
For example, Chang et al. (2024) measures $S(\omega)$ using the perplexity measure.
Turning to Goyal et al. (2024), we can define a notion of the declining value of data as the number of training epochs $k$ increases (and hence the amount of repeated data increases). We can define an additive term in the data deficiency as follows:
$\mu_{3}\frac{k}{\tau}$
where $k$ is the number of epochs. With this form, for small number of epochs, the loss decreases, but as the number of epochs increases, the loss begins to grow exponentially, and then levels out, with the rate determined by a “decay parameter” $\tau$.
Appendix E Additional Validations and Practical Guide
In this section, we present additional experiments validating our noise model and scaling law stability, along with a practical guide for estimating $Q$.
E.1 Validation of Synthetic Noise Model
To validate that our synthetic noise injection serves as a meaningful proxy for real-world data degradation, we measured the semantic drift of the corrupted data. We computed the average cosine similarity between the embeddings of clean samples and their synthetically corrupted counterparts using a pre-trained reference model. As shown in Figure 4 there is a strictly monotonic, near-linear increase in embedding similarity as the synthetic quality increase (noise level decrease). This confirms that our parameter $Q$ effectively captures the loss of semantic information.
Figure 4: Semantic drift analysis: Average cosine similarity between embeddings of clean and corrupted samples at various quality levels.
E.2 Sensitivity to Learning Rate
We performed a sensitivity analysis by retraining models at 5 different learning rates (1e-5 to 5e-3) across 3 dataset scales and 3 noise levels. As shown in Table 9 and Figure 5, the estimated $\gamma$ is highly stable ($\sim 0.39$) for practical learning rates near the optimal 1e-3 used in our main experiments.
Table 9: Sensitivity of the estimated quality exponent $\gamma$ to learning rate.
Learning Rate Estimated $\gamma$
1e-5 0.037 (Undertrained)
3e-4 0.394
1e-3 (Ours) 0.396
5e-3 0.177 (Unstable)
(a) $\gamma$ vs learning rate
(b) Test loss vs learning rate
Figure 5: Sensitivity analysis of learning rate: (a) Estimated $\gamma$ across different learning rates, (b) Learning rate vs test loss for different batch sizes.
E.3 Practitioner’s Guide to Estimating Q
We propose the following recipe for estimating $Q$ in practice:
  1. 1.
    Establish a baseline: Identify a small, high-quality reference subset (e.g., manually curated text or a trusted domain corpus) to represent $Q\approx 1$.
  2. 2.
    Select quality proxies: Choose measurable proxies for “deficiency” relevant to the domain.
  3. 3.
    Compute deficiency ($\Delta$): Construct a composite deficiency score $\Delta=\sum_{j}\mu_{j}\Delta_{j}$, where $\mu_{j}$ are weights determined by importance or preliminary ablation.
  4. 4.
    Map to $Q$: Calculate $Q=e^{-\Delta}$. This maps the unbounded deficiency metric to the $(0,1]$ interval.
  5. 5.
    Sanity check: Verify that $Q$ correlates with downstream performance or semantic drift (as in our embedding experiment) on a small validation set.
  6. 6.
    Fit scaling law: Use the estimated $Q$ values for different data subsets to fit the parameters ($B,\beta,\gamma$) of the scaling law $L\approx B/(D^{\beta}Q^{\gamma})+E$ (assuming $N$ is fixed or large enough), using the robust Huber loss method described in the paper.